The diagonal brace on a rectangular gate forms an angle with the horizontal side with degree x and an angle with the vertical side with degree measure 2x-3. Find x.
How do I dolve this.

Answer by venugopalramana(3286)

The diagonal brace on a rectangular gate forms an angle with the horizontal side with degree x and an angle with the vertical side with degree measure 2x-3. Find x.
How do I dolve this.
TOTAL ANGLE =90
X+2X-3=90
3X=90+3=93
X=31